OpenStudy (anonymous):

The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (gorv):

what is the formula for area of rectangle ???

OpenStudy (anonymous):

A = wl

OpenStudy (gorv):

let width=x

OpenStudy (gorv):

A=240=x*l cal u tell me value of l in terms of x??

OpenStudy (anonymous):

I really don't understand this stuff :(

OpenStudy (anonymous):

id say 240=w(w+8)

OpenStudy (gorv):

divide both side by x

OpenStudy (anonymous):

Where does the 8 come in?

OpenStudy (anonymous):

From*

OpenStudy (anonymous):

which is actually w^2+8w=240

OpenStudy (gorv):

@countrygirl012 plz follow my approach

OpenStudy (anonymous):

Okay.. @gorv

OpenStudy (yanasidlinskiy):

Using as what you've said above, the area is 240. Now, all you have to do is plug in the numbers. Width = w and l = length. So your equation would look something like this: A= 240 = 5*3 8 is wrong, because you're supposed to be multiplying the length and the width not adding. The formula is: A=l*w

OpenStudy (gorv):

A=l*w 240=l*w

OpenStudy (anonymous):

Thank you!! @YanaSidlinskiy May I ask what I do for Part B??

OpenStudy (yanasidlinskiy):

No problem. For B, all they're asking for is the steps on how you solve it.

OpenStudy (yanasidlinskiy):

So just solve it until you cannot simplify anymore further.

OpenStudy (gorv):

w=240/L

OpenStudy (anonymous):

So because it asking for me to find the width... I would change the equation to look something like this... W= 240/L right?

OpenStudy (gorv):

If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. which means width=240/L+3 now and length= L-5

OpenStudy (gorv):

in square length=width

OpenStudy (anonymous):

I am really confused.. .-.

OpenStudy (gorv):

\[L-5=\frac{ 240 }{ L}+3\]

OpenStudy (gorv):

now solve this

OpenStudy (anonymous):

Is that how i find the width

OpenStudy (gorv):

yeah

OpenStudy (gorv):

5 and 3 are not width

OpenStudy (gorv):

5*3= 15 not 240

OpenStudy (anonymous):

-12, 20???

OpenStudy (anonymous):

I solved for l though ...

OpenStudy (gorv):

yes

OpenStudy (gorv):

but length or width can not be negative

OpenStudy (gorv):

so we left with only 20

OpenStudy (gorv):

L=20

OpenStudy (gorv):

\[w=\frac{ 240 }{ L }\]

OpenStudy (anonymous):

Okay so do I plug it in?

OpenStudy (gorv):

plug in 20 in place of L and you will get value of width

OpenStudy (anonymous):

\[w = \frac{ 240 }{ 20 }\]

OpenStudy (gorv):

solve it further and tell me what you got ???

OpenStudy (anonymous):

12?

OpenStudy (gorv):

yep you got it right :)

OpenStudy (gorv):

width = 12

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (gorv):

u r welcome :)